# Group-On: Boosting One-Shot Segmentation with Supportive Query

### v1

@@ 3.4 Group-On Voting Module
5-support HSNet5 [20] 45.90 53.00 51.80 47.10 49.50 72.40 71.80 74.40 67.00 68.30 70.40 80.60
ASNet5 [1] 48.00 52.10 49.70 48.20 49.50 72.70 73.10 75.60 65.70 69.90 71.10 81.00
1-support SSP [32] 39.10 45.10 42.70 41.20 42.00 - 63.20 70.40 68.50 56.30 64.60 -
DCAMA [14] 41.50 46.20 45.20 41.30 43.50 69.90 65.40 71.40 63.20 58.30 64.60 77.60
VAT [21] 39.50 44.40 46.10 40.40 42.60 - 68.40 72.50 64.80 64.20 67.50 78.80
QCLNet [39] 40.00 45.50 45.10 43.60 43.60 - 67.90 72.50 64.30 63.40 67.00 -
SCCAN [41] 42.60 51.40 50.00 48.80 48.20 69.70 70.00 73.90 66.80 61.70 68.30 78.50
HSNet [20] 37.20 44.10 42.40 41.30 41.20 69.10 67.30 72.30 62.00 63.10 66.20 77.60
ASNet [1] 41.80 45.40 43.20 41.90 43.10 69.40 69.00 73.10 62.00 63.60 66.90 78.00
HSNet + Group-On 44.87 53.44 48.59 47.73 48.66 (7.46 $\uparrow$ ) 73.43 (4.33 $\uparrow$ ) 72.57 74.99 63.53 66.82 69.48 (3.28 $\uparrow$ ) 80.38 (2.78 $\uparrow$ )
ASNet + Group-On 48.08 54.18 51.64 51.34 51.31 (8.21 $\uparrow$ ) 73.46 (4.06 $\uparrow$ ) 73.83 75.51 62.48 69.25 70.27 (3.37 $\uparrow$ ) 80.24 (2.24 $\uparrow$ )
Table 2: Segmentation performance on FSS-1000 [7] in the 1-support $C$-query setting. HSNet5 and ASNet5 represent the results of the baseline models HSNet and ASNet in the conventional 5-support setting, while the rest results are all in the 1-support setting. ASNet: Our reimplementation based on official codes. We mark the best results in bold.
Setting Method ResNet-50 ResNet-101
mIoU FB-IoU mIoU FB-IoU
5-support HSNet5 [20] 87.50 - 88.50 -
ASNet5 [1] 87.28 92.07 88.15 92.65
1-support SSP [32] 87.30 - 88.60 -
Anno-free FSS [42] 85.70 - - -
DifFSS [43] 88.40 - - -
HSNet [20] 85.50 - 86.50 -
ASNet [1] 85.96 91.04 86.99 91.67
HSNet + Group-On 87.99 92.62 88.24 92.65
ASNet + Group-On 88.21 92.63 88.86 93.04
4 Experiments
Datasets.
@@ 4 Experiments
We evaluate our Group-On approach on three standard benchmarks: COCO-$20^{i}$ [6], PASCAL-$5^{i}$ [5], and FSS-1000 [7]. Following previous work [20, 44], the experiments on COCO-$20^{i}$ and PASCAL-$5^{i}$ are conducted in a cross-validation manner over four folds, with 20 and 5 classes per fold, respectively. For each fold, 1000 episodes are sampled for evaluation, while the rest of the data is used for training. The FSS-1000 dataset containing 1000 classes is divided into 520, 240, and 240 classes for training, validation, and test.
Table 3: Segmentation Performances on 4 folds of the PASCAL-5i [5] dataset with CLIP [45] backbone under the 1-support $C$-query setting. HSNet, HSNet5: Our reimplementation of the baseline model HSNet based on official codes.
Setting Method Fold-0 Fold-1 Fold-2 Fold-3 Mean FB-IoU
5-support HSNet5 77.07 82.19 73.43 76.47 77.29 86.49
1-support HSNet 72.53 76.38 68.58 69.58 71.77 82.59
HSNet + Group-On 77.13 80.47 68.82 74.60 75.26 (3.69 $\uparrow$ ) 84.67 (2.08 $\uparrow$ )
Table 4: FPS on 4 folds of the PASCAL-5i [5] dataset with ResNet50 [37] under the 1-support $C$-query setting. HSNet5 represents the results of the baseline model HSNet in the conventional 5-support setting, while the rest results are all in the 1-support setting.
Setting Method Fold-0 Fold-1 Fold-2 Fold-3 Mean
5-support HSNet5 [20] 4.90 4.77 4.77 4.90 4.83
1-support HSNet [20] 23.33 22.00 21.67 24.00 22.75
HSNet + Group-On 6.51 6.26 6.34 6.88 6.50
Implementation Details.
@@
We incorporate Group-On into two baseline models, HSNet [20] and ASNet [1], denoted as HSNet+Group-On and ASNet+Group-On and they are built upon the PyTorch [46] framework. We use 2 query images for training and 5 for testing, and add the step of alignment regularization [23] to help the training. The backbones ResNet50 [37] and ResNet101 [37], which have been pre-trained on ImageNet [47], are fixed to extract features (features from conv3  x to conv5  x before the ReLU [48] activation of each layer are stacked to form the deep features). In addition, we also experiment with the public pretrained model ViT-L/14@336px of CLIP [45], to evaluate the generalization of our approach on the non-convolutional backbone. In the Group-On Voting module, we select features from the 9th layer as input. The default settings for the optimizer and learning rate of the two baselines remain unchanged. For the batch size, we use 64 for HSNet+Group-On and 48 for ASNet+Group-On. Model training is set to halt when it reaches the maximum 500th epoch. Note that our experiments demonstrate a wide coverage spanning across multiple datasets to various backbones in order to validate the effectiveness and generalizability of our Group-On approach, and they are done on four NVIDIA A100 Tensor Core GPUs for over two months.
Evaluation Metrics.
@@
We evaluate the performance with two widely-adopted metrics: mean intersection over union (mIoU) and foreground-background IoU (FB-IoU) [20, 44]. They are computed as:
$\text{mIoU}=\frac{1}{n}\sum_{i=1}^{n}\text{IoU}_{i},\vspace{-0.2cm}$ (8)
where n is the number of test samples.
@@
where f means foreground and b means background without regard to the object classes.
@@
4.1 Comparison with State-of-the-Art Methods
4.1.1 Quantitative Results
We compare our results with recent methods [14, 32, 21, 4, 39, 41, 43, 42]. Unless otherwise specified, the reported results of the known methods are all from the original papers. The results in Tab. 1 (left) are on the COCO-20i dataset and provide the most noticeable gains. Particularly, HSNet+Group-On and ASNet+Group-On with ResNet101 obtain significant improvements of 7.46% and 8.21% on mIoU compared to the counterpart, respectively. The main reason of the superior improvement is the more severe intra-class variation issue derived by the larger data volume in COCO-20i dataset. Tab. 1 (right) shows the results on the PASCAL-5i dataset. Our approach delivers consistent improvement (average over 3% in mIoU and 2% in FB-IoU). Tab. 2 represents the results of our approach in the 1-shot setting on the FSS-1000 dataset, which are all competitive with the counterparts in the 5-shot setting. HSNet+Group-On delivers a gain of 2.49% with ResNet50 and 1.74% with ResNet101 in mIoU. Furthermore, Tab. 3 shows the results of our additional experiments with CLIP on the PASCAL-5i dataset. Given both time and GPU memory considerations, we experiment with one baseline model on one dataset for reference. When using the strong backbone CLIP as the feature extractor, HSNet+Group-On still outperforms the baseline model HSNet by a margin of 3.69% and 2.08% in mIoU and FB-IoU respectively in the 1-shot setting. Such improvements further underscore the capability of our approach to effectively harness the diversity among queries and demonstrate its broad applicability to any backbone, even the most cutting-edge one.
The inference efficiency of our approach is measured by frames per second (FPS). The experiments are carried out on PASCAL-5i dataset, and we average the results from 5 runs. Regarding the 1-shot setting in Tab. 4, we sacrifice some time cost of inference for better performances. Since our performance enhancement is considerably big and even competitive to the counterparts in the 5-shot setting, its efficiency is acceptable. In comparison to the 5-shot setting, our method is not only more efficient in terms of inference time alone, but also reduces the required manual annotation time for the few-shot setting.
4.1.2 Qualitative Results
To intuitively display our approach, we visualize some results containing challenging intra-class variation under the one-shot setting in Fig. 3. The first row is support images with their masks in blue, while the second row represents the groundtruth. And the following two rows display predictions of the baseline and our approach, respectively. All the images in each column belong to the same category. One can see from our results (the 4th row) that if objects in the support image differ greatly from those in the query images, Group-On captures the semantic essence of the objects and segments more accurately with the reciprocal knowledge help of the multiple query images.
4.2 Ablation Study and Analysis
Extensive ablation experiments from multiple aspects are conducted to provide more insights into our approach, using the PASCAL-5i [5] dataset and ResNet-50 [37] backbone.
4.2.1 Analysis of Attention Correlations for Group-On Voting
We conduct experiments to validate that cross-attention correlations between the host and supportive query reflect deep relations of image pairs, which guide the assignment of voting weights. We use an indicator $\psi$ to assess the scene differences of two images [26]. Defined by Eq. (10), $\psi$ is the Frobenius norm evaluated on the difference of two Gram matrices [26] $\textbf{{G}}^{h}$ and $\textbf{{G}}^{s}$:
$\psi=-\|\textbf{{G}}^{h}-\textbf{{G}}^{s}\|_{F},$ (10)
where $h$ and $s$ denote the host query and supportive query, respectively. More complete calculation details are in Appendix A. We randomly sample 100 image pairs from each category and calculate the average results of correlation coefficients for 10 runs. As shown in Fig. 5, the Pearson coefficient [49] and Spearman coefficient [50] between $\psi$ and cross-attention correlations are mostly above 0.5 (which happens for 17 out of 20 for Pearson and 15 out of 20 for Spearman). Such phenomenon indicates a strong connection between the scene difference of images (represented by $\psi$) and cross-attention correlations (used in Group-On Voting), and thus verifies the effectiveness of the Group-On Voting module.
Figure 4: Analysis of attention correlations for Group-On Voting on PASCAL-5i [5] dataset. We randomly sample many image pairs from totally 20 categories [36] and calculate correlation coefficients between attention correlations and scene differences of image pairs.
Figure 5: Ablation study on the Group-On Voting weights on the PASCAL-5i dataset. att: cross-attention voting weight scheme. arith: arithmetic mean scheme. In the legend, the former one is for training while the latter for testing. We investigate the best combination of att and arith, and (att, arith) is the optimal.
Figure 6: Comparison of the impact of query amounts in training and testing, performed on the fold-0 of PASCAL-5i. In the left, the best option of query amount for training is 2, given its slight advantage. In the right, with the growth of query amount, the inference performance climbs up and reaches a stable fluctuation. We select 5 queries for testing.
4.2.2 Ablation Study on Voting Weights
We design ablation study to examine how to assign voting weights effectively. The cross-attention voting weight and arithmetic mean schemes are denoted as att and arith respectively in Fig. 5. In the legend, the former one is for training while the latter for testing. The performance gaps between (att, arith) and the others on four folds confirm that our proposed option is the best. That is, the Group-On Voting module assigns the voting weights by cross-attention correlations for training, and by arithmetic mean schemes for testing.
4.2.3 Impacts of Query Amounts
We perform comparative experiments using varying numbers of query images during training and testing. Fig. 6 (top) shows that except for the baseline setting (i.e., $amount=1$), the results of using the other query amounts rise to a similar level and the option of $amount=2$ wins narrowly. Considering more calculation afflictions as the number of query images increases, we fix the query number to 2 for training here. In Fig. 6 (bottom), with the growth of query number for testing, the performance gradually increases and eventually reaches a stable fluctuation. Due to space limitations, we represent relations between more query amounts and inference performances in Appendix B. For the sake of alignment with the common 5-shot setting and less computation, here we select 5 query images for testing per episode. More exactly, Group-On tests query images in batches of 5 for complementary help of one another. In general, the performance can be improved so long as more query images than one are used, and one may select the most favorable query amount.
5 Conclusions
We proposed a simple and novel approach called Group-On to address the intra-class variation challenge in ONE-shot segmentation. The core innovation of Group-On is to leverage mutual information benefits of query images in a group for more representation multiplicities in a category. Experiments on three standard benchmarks showed significant performance enhancement and the effectiveness of our approach, which achieved comparable results with the counterparts under the 5-shot setting. In essence, we boosted one-shot segmentation from the data-centric perspective instead of the model architecture design and data labeling. We hold the expectation that our work will shed light on future research in addressing the intra-class variation issue.
Appendix A Analysis of Attention Correlations for Group-On Voting

### v2

@@ II-B Group-On Process under One-Shot Setting
5-support HSNet5 [6] 45.9 53.0 51.8 47.1 49.5 72.4 71.8 74.4 67.0 68.3 70.4 80.6
ASNet5 [1] 48.0 52.1 49.7 48.2 49.5 72.7 73.1 75.6 65.7 69.9 71.1 81.0
1-support QCLNet [13] 40.0 45.5 45.1 43.6 43.6 - 67.9 72.5 64.3 63.4 67.0 -
SCCAN [14] 42.6 51.4 50.0 48.8 48.2 69.7 70.0 73.9 66.8 61.7 68.3 78.5
HSNet [6] 37.2 44.1 42.4 41.3 41.2 69.1 67.3 72.3 62.0 63.1 66.2 77.6
ASNet [1] 41.8 45.4 43.2 41.9 43.1 69.4 69.0 73.1 62.0 63.6 66.9 78.0
Group-On (HS) 44.9 53.4 48.6 47.7 48.7 (7.5 $\uparrow$ ) 73.4 (4.3 $\uparrow$ ) 72.6 75.0 63.5 66.8 69.5 (3.3 $\uparrow$ ) 80.4 (2.8 $\uparrow$ )
Group-On (AS) 48.1 54.2 51.6 51.3 51.3 (8.2 $\uparrow$ ) 73.5 (4.1 $\uparrow$ ) 73.8 75.5 62.5 69.3 70.3 (3.4 $\uparrow$ ) 80.2 (2.2 $\uparrow$ )
Model performances on the CLIP backbone
1-support PGMA-Net [15] - - - - - - 74.0 81.9 66.8 73.7 74.1 82.1
HSNet [6] 53.4 60.5 58.3 57.3 57.4 75.0 72.5 76.4 68.6 69.6 71.8 82.6
ASNet [1] 53.6 60.8 57.8 56.4 57.1 75.1 74.4 77.8 68.3 73.2 73.4 82.8
Group-On (HS) 54.5 62.8 58.5 57.5 58.3 (0.9 $\uparrow$ ) 78.1 (3.0 $\uparrow$ ) 77.1 80.5 68.8 74.6 75.3 (3.5 $\uparrow$ ) 84.7 (2.1 $\uparrow$ )
Group-On (AS) 55.0 61.8 59.0 57.9 58.4 (1.3 $\uparrow$ ) 78.2 (3.1 $\uparrow$ ) 77.9 79.7 68.3 76.9 75.7 (2.3 $\uparrow$ ) 84.2 (1.4 $\uparrow$ )
II-C Mixture of Mask Experts Module
For advanced organization and utilization of candidate masks, we propose a novel Mixture of Mask Experts (MoME) module inspired by MoE. Since candidate masks are predicted based on supportive queries/support image that contain intra-class variation, there is an urgent need for multiple mask experts proficient in different aspects to handle the candidate masks effectively. Since the support image and supportive queries function similarly in this module, they are collectively referred to as supportive query data below for simplicity.
As illustrated in Fig. 2, a scene-driven router evaluates and scores expert decisions based on query image scenes, with each mask expert responsible for one candidate and working together to overcome the intra-class variation challenge. Firstly, a scene-driven router (Router) evaluates the contribution of each expert based on scene differences between host and supportive queries, generating router weights $W$ through cross-attention:
$W=\texttt{Router}(\textit{F}^{\text{host}},\textit{F}^{\text{supp}}_{1},\textit{F}^{\text{supp}}_{2},\ldots,\textit{F}^{\text{supp}}_{N_{q}}),$ (4)
where $W=[w_{1},\ldots,w_{i},\ldots,w_{N_{q}}]$ and $w_{i}$ is the router weight of $i_{th}$ candidate mask. Steered by the router, secondly we form $N_{q}$ decisions by feeding candidate masks into an ensemble of mask experts (Expert), each of which comprises a convolutional layer with different weights. Eventually, we obtain the elegant final result $\hat{\textit{M}}^{\text{host}}$ through a weighted sum of mask expert decisions:
$\hat{\textit{M}}^{\text{host}}=\sum_{i=1}^{N_{q}}\texttt{Expert}(\textit{{C}}_{i})*w_{i}.$ (5)
With a flexible number of mask experts (i.e., query amount $N_{q}$), MoME module facilitates more efficient integration of candidate masks and significantly expands the model capability to handle different types of images. Furthermore, such module provides robust technical assurance for mutual benefits of multiple queries in Group-On, alleviating severe intra-class variation issue.
III Experiments
Datasets and Metrics
@@ III Experiments
We evaluate Group-On on three standard benchmarks: COCO-$20^{i}$ [5], PASCAL-$5^{i}$ [9], and FSS-1000 [16]. PASCAL-$5^{i}$ and COCO-$20^{i}$ contain 20 and 80 object classes, respectively. And experiments on them are conducted in a cross-validation manner over four folds. For each fold, 1000 episodes are sampled for evaluation, while samples from the rest three folds are used for training. The FSS-1000 dataset containing 1000 classes is divided into 520, 240, and 240 classes for training, validation, and test. We use two widely-adopted metrics [6]: mean intersection over union (mIoU) and foreground-background IoU (FB-IoU).
Figure 3: Ablation study of MoME module.
Figure 4: Interpretability of the router in MoME module.
Implementation Details
We incorporate Group-On into two baseline models using the PyTorch framework, HSNet [6] and ASNet [1], denoted as Group-On (HS) and Group-On (AS). We use 2 query images for training and 5 for testing, and add alignment regularization [17] to facilitate training. For backbone networks, we apply ResNet50 and ResNet101 [18] and they are fixed in training. In addition, we experiment with the public pretrained model ViT-L/14@336px of CLIP [19], to evaluate the capability of our approach on the modern strong backbone. More details are in Appendix.
III-A Comparison with State-of-the-Art Methods
In this subsection, we compare Group-On with multiple recent methods. The results in Tab. I (left) are on the COCO-20i dataset and provide the most noticeable gains. Particularly, Group-On (HS) and Group-On (AS) with ResNet101 obtain significant improvements of 7.5% and 8.2% on mIoU compared to the counterpart, respectively. Tab. I (right) shows the results on the PASCAL-5i dataset. Our approach delivers consistent improvement (average over 3% in mIoU). Tab. II reports results of our approach in the 1-shot setting on the FSS-1000 dataset, which are all competitive with the counterparts in the 5-shot setting.
Furthermore, the bottom of Tab. I shows results of additional experiments with CLIP on the two datasets. When using the strong backbone CLIP as the feature extractor, our Group-On approach still outperforms baseline models and recent methods under the 1-shot setting. In particular, Group-On (HS) obtain the improvements of 3.5% in mIoU on the PASCAL-20i dataset. Such improvements further underscore the capability of our approach to effectively harness the diversity among queries and demonstrate its broad applicability to any backbone, even the most cutting-edge one. Furthermore, please see visualization results in Appendix.
III-B Ablation Study and Analysis
We conduct extensive ablation experiments using the PASCAL-5i [9] dataset and ResNet-50 [18] backbone.
Ablation Study on MoME module
How to effectively integrate candidate masks of the host query is crucial for fully unlocking the potential of supportive queries in mutual assistance. Thus we compare three alternative choices of candidate mask integration in Fig. 4. arith: arithmetic mean scheme by calculating the average of all masks. union: foreground union scheme by taking the union of foregrounds in all masks. Obviously, the performance gaps between MoME and the others on four folds confirm the superiority of MoME module.
TABLE II: Performance on FSS-1000 [16] in 1-support $N_{q}$-query setting. ASNet5: results in the conventional 5-support setting. ASNet: Our reimplementation.
Setting Method mIoU FB-IoU
Model performances with the ResNet-50 backbone
5-support ASNet5 [1] 87.3 92.1
1-support DifFSS [20] 88.4 -
ASNet [1] 86.0 91.0
Group-On (AS) 88.2 92.6
Model performances with the ResNet-101 backbone
5-support ASNet5 [1] 88.1 92.6
1-support ASNet [1] 87.0 91.7
Group-On (AS) 88.9 93.1
Interpretability of Scene-driven Router
We aim to enhance interpretability of the scene-driven router that guides the mixture of mask expert decisions in MoME module. As a widely-used metric in the field of style transfer, $\psi$ measures scene differences between two images (Calculation details in Appendix) and previous practice [21] has verified its effectiveness. In this experiment, we sample 100 random image pairs from each category to compute the tried-and-true factor $\psi$ and router weights, and then calculate correlation coefficients between these two factors. In Fig. 4, Pearson and Spearman coefficients between $\psi$ and router weights are mostly above 0.5 (17/20 for Pearson and 15/20 for Spearman). Such phenomenon signals that $\psi$ and router weights are positively related, and the scene-driven router guides expert decisions based on scene differences between host and supportive queries. Some visualization results are in Fig. 5.
Impacts of Query Amounts
The investigation of varying query amounts (i.e., mask expert amount) is in Fig. 6. As the number of query images increases, the inference performance first climbs up and then remains relatively stable. However, the inference time per query manifests an sustainable growth. Such trends indicate a marginal performance improvement at the cost of significant computational resources after reaching a point of stable fluctuation. Therefore, one should select the most favorable query amount that balances both efficiency and performance, and in this paper we set the query amount to 2.
Figure 5: Visualization results of different supportive queries and their corresponding router weights.
Figure 6: Impacts of varying number of queries. We set the query amount to 2 given the balance of efficiency and performance.
IV Related Work
We provide more detailed related works in this section. Few-shot segmentation is to make segmentation in the low-data regime. There are numerous methods, most of which investigated interactions between support and query images while the others focused on details inside either the support or query data.
Interactions between Support and Queries.
Many studies [4, 22] sought to design high-capacity interaction modules between the support and query data. It was common to expand features in multiple scales and layers [23] to enhance in-depth interplay. Based on the Transformer architecture, CyCTR [24], DCAMA [25], and HDMNet [26] developed cross-attention operations while some others [27, 3, 28, 29, 30] did not. Besides, there were some attempts to generate high-dimensional feature correlation tensors, such as ASNet [1], VAT [7]. However, they tended to suffer from complex and expensive computations.
Support Exploration.
Due to the scarcity of reference information, some studies took advantage of the support data in various perspectives (e.g., background and boundaries). PANet [17] proposed to take the predicted query data as pseudo references to segment support images in turn. As the knowledge of objects from base classes in the background can be ample, it was investigated [31, 21] to use such knowledge to assist novel class predictions. HM [32] and MSI [33] employed two masking techniques to segment more accurate target object boundaries. Recently, MIANet [34] introduced supplementary category labels to leverage the general semantic information in a multi-modal manner.
Query Exploration.
For query exploration, DPNet [35] added pseudo-prototypes extracted from the query foreground to the original support prototypes for refinement. SSP [36] suggested that the locality of query masks, which were predicted in high confidence, could aid the support data to segment the leftover low-confidence areas. The known methods [37] on query exploration mainly aimed to make full use of what is inside a single query image to guide the segmentation. In contrast, our Group-On approach exploits directly increasing the query amount to enrich semantic knowledge in an intuitive way. Group-On is more capable and simpler, and can be incorporated seamlessly with plentiful FSS baselines.
V Conclusions
We proposed a novel and effective approach called Group-On to address the intra-class variation challenge in ONE-shot segmentation. The core innovation of Group-On is to leverage mutual information benefits of query images in a group for more representation multiplicities in a category. To make the most of multiple query images, we construct the MoME module containing a scene-driven router and a flexible number of mask experts that jointly make comprehensive decisions. Experiments on three standard benchmarks showed significant performance enhancement and the effectiveness of our approach, which achieved comparable results with the counterparts under the 5-shot setting. In essence, we boosted one-shot segmentation from the data-centric perspective instead of the model architecture design and data labeling. We expect that our work will shed light on future research in addressing the intra-class variation issue.
Acknowledgment
This research was partially supported by National Natural Science Foundation of China (No. 12326612), Zhejiang Key R&D Program of China (No. 2023C03053), Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence, State Key Laboratory of Transvascular Implantation Devices (NO. SKLTID2024003), and the Transvascular lmplantation Devices Research Institute (TIDRI).
Appendix A Experiments
A-A Implementation Details.
@@ Appendix A Experiments
We provide more implementation details in this section. For backbone networks, we do not include VGG-16 [38] due to its inferior performances compared with the ResNet family in the previous works, and instead we apply ResNet50 [18] and ResNet101 [18], which are frozen to prevent them from learning class-specific representations of the training data [6]. They have been pre-trained on ImageNet [39] and are fixed to extract features (features from conv3  x to conv5  x before the ReLU [40] activation of each layer are stacked to form the deep features). In the MoME module, we select features from the 9th layer as input. The default settings for the optimizer and learning rate of the two baselines remain unchanged. Model training set to halt when it reaches the maximum 500th epoch. Note that our experiments demonstrate a wide coverage spanning across multiple datasets to various backbones in order to validate the effectiveness and generalizability of our Group-On approach, and they are done on four NVIDIA A100 Tensor Core GPUs. For evaluation metrics, mIoU averages over IoU values of all classes in the testing data to avoid the bias to the specific categories, while FB-IoU ignores object classes and computes average of foreground and background IoUs. As mIoU better reflects model generalization capability and prediction quality than FB-IoU does [6], we mainly focus on mIoU in our experiments.
A-B Evaluation Metrics.
@@
We evaluate the performance with two widely-adopted metrics [6, 41]: mean intersection over union (mIoU) and foreground-background IoU (FB-IoU), as:
$\text{mIoU}=\frac{1}{n}\sum_{i=1}^{n}\text{IoU}_{i},$ (6)
@@
where n is the number of test samples, and f means foreground and b means background without regard to the object classes.
A-C Analysis of Datasets
Due to more severe intra-class variation in COCO-20i, Group-On achieves greater performance improvements in COCO-20i than that in PASCAL-5i. In this section, we conduct analysis about the intra-class variation in these two datasets. There are overlapped 20 categories in COCO-20i and PASCAL-5i, and we utilize the Affinity Propagation clustering algorithm within each category to cluster images based on image similarity. As represented in Tab. III, the average amount of clusters in COCO-20i is almost over four times than that in PASCAL-5i. This further reflects the more severe intra-class variation issue in COCO-20i dataset from an intuitive perspective.
TABLE III: The number of clusters in each category of PASCAL-5i and COCO-20i datasets.
category aeroplane bicycle bird boat bottle bus car cat chair cow diningtable dog horse motorbike person pottedplant sheep sofa train tvmonitor avg
pascal 15 14 20 11 12 12 28 30 30 7 20 37 13 12 84 11 7 14 14 17 20.4
coco 59 51 42 42 82 67 127 68 159 33 148 63 57 64 514 64 35 88 62 96 96.1
@@ A-C Analysis of Datasets
A-D Interpretability of Router Weights in MoME